### v1

# HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models

Abstract
The ubiquitousness of social media has led to the need for reliable and efficient detection of offensive content to limit harmful effects. This has led to a proliferation of datasets and models related to detecting offensive content. While sophisticated models have attained strong performance on individual datasets, these models often do not generalize due to differences between how “offensive content” is conceptualized, and the resulting differences in how these datasets are labeled. In this paper, we introduce HateCOT, a dataset of 52,000 samples drawn from diverse existing sources with explanations generated by GPT-3.5-Turbo and human-curated. We show that pre-training models for the detection of offensive content on HateCOT significantly boots open-sourced Language Models on three benchmark datasets in both zero and few-shot settings, despite differences in domain and task. We further find that HateCOT enables effective K-shot fine-tuning in the low-resource settings. Our repository is available at https://github.com/hnghiem-usc/hatecot.
1 Introduction
As social media has become an indispensable component of modern discourse, this channel of communication has also amplified the propagation of offensive content. Speech aimed to promote hateful sentiments thrive on such platforms, leading to real and significant harm on their audience Giachanou and Rosso (2020); Saha et al. (2019). However, “offensive content” is an essentially contested construct, and exactly what is and is not allowed varies by platform. In research, different approaches analyze semantically similar but still distinct concepts: Cyber-bullying, Toxicity, Sexist, Racist, Hate etc. Poletto et al. (2021); Fortuna et al. (2021); Nghiem et al. (2024), making visible this contestedness.
Compounding the challenge, reliable detection of offensive content requires significant data resources. Sophisticated models tend to be data-hungry, and the process of curating a dataset tailored to specific use cases can be costly, time-consuming, and emotionally challenging for annotators Founta et al. (2018); Toraman et al. (2022). The typical pipeline consists of collecting samples based on topic-relevant key words, then recruiting either crowdworkers or experts to annotate data before developing classification models Paullada et al. (2021). Each step incurs investment and may inject subtle idiosyncrasies to the downstream dataset, further limiting potential transferable usefulness to related tasks Fortuna et al. (2021).
In practical settings, users may desire transparency from social media platforms. Therefore, the ability to provide human understandable justification based on platform-specific policy becomes an attractive feature for content moderation. Nevertheless, current techniques often still fail to offer intuitive explanatory signals, especially to end users Yadav et al. (2023); Babaeianjelodar et al. (2022); Ibrahim et al. (2022).
In this work, we attempt to simultaneously reduce the cost of data curation, enhance cross-dataset generalization, and address the necessity of explainable decisions for offensive content detection. Specifically, our main contributions are:
  1. 1.
    We curate and release HateCOT (Hate-related Chains-of-Thought) a dataset of $52k$ samples consisting of input text, a hate speech label, and an explanation of that label. We collect this data by merging eight datasets with explanations created by using GPT-3.5-Turbo to augment human annotations.
  2. 2.
    We demonstrate the efficacy of using HateCOT as a pre-training corpus to enable finetuning on a target domain and task. We show that open-sourced LLMs can effectively leverage definitions to adapt to new tasks using zero-shot and few-shot settings via finetuning and in-context learning across 3 datasets.
  3. 3.
    We assess the quality of explanations generated by our finetuned models with respect to the criteria described in their corresponding dataset’s definitions. These insights showcase LLM-generated explanations as a means to enhance transparency in content moderation.
2 Related Works
2.1 Offensive Speech Detection
Offensive speech detection has attracted considerable interest from the research community. Earlier approach typically investigated coarse-grained labels (e.g. Hate vs Not Hate) while subsequent efforts explored more diverse facets of offensive speech at higher granularity Founta and Specia (2021); Poletto et al. (2021); Vidgen and Derczynski (2020). Increasingly more advanced models emerged over time with the diversity of datasets.
Cross-domain generalization, however, still remains a relevant challenge in the area. Fortuna et al. (2021) found empirically that cross-dataset transference is highly dependent on semantic similarity between their label spaces. Recent works have pre-trained Transformer-based models, such as HateBERT and fBERT, on specialized corpora to enhance generalization to various levels of success Caselli et al. (2021); Sarkar et al. (2021).
2.2 LLMs in Offensive Speech Classification
Zampieri et al. (2023) assessed a range of open-sourced LLMs on zero-shot prompting on the OffensEval task and found their performance trailing by a wide margin behind trained existing fine-tuned BERT-based systems. Chiu et al. (2021) and Han and Tang (2022) used the proprietary GPT-3 on a set of different datasets and noted that informative contexts and examples could boost the model’s performance to competitive levels on a different set of data. Similarly, Roy et al. (2023) found that adding explanation to pipeline could result in 10 to 20% boost in performance of LLMs over baselines. (Yang et al., 2023)’s study found that training LLMs with step-by-step reasoning grounded by annotations could improve predictive power.
Pre-trained language models have exhibited remarkable ability in text generation Celikyilmaz et al. (2020). Recent scaled-up LLMs such as the GPT-3 and later models are capable of astoundingly fluent, convincing and knowledge-infused outputs Zhang et al. (2023). LLMs with hundreds or more parameters even exhibit reasoning capabilities Wei et al. (2022, 2021), leading to a flurry of research on prompting techniques to harness their prowess, such as Chain-of-thought (COT), Tree-of-thought etc. (Yao et al., 2023; Diao et al., 2023). An interesting line of research leverages LLMs to efficiently generate high volumes of synthetic data for tasks with training resource is scarce Puri et al. (2020); Bao et al. (2023); Whitehouse et al. (2023). We build upon these works to construct a dataset that can induce smaller LLMs to efficiently adapt to new categories of offensive content by leveraging their provided definitions.
3 Building HateCOT
We first describe the process to identify the candidate datasets from literature, and the procedure to obtain annotation-guided explanations from these samples (Section 3.1). We then perform a set of validation experiments to optimize the data’s parameters for downstream tasks before augmenting our corpus to its eventual size (Section 3.2).
3.1 Data Selection
Datasets for Training.
We use the following criteria to filter existing corpora related to offensive speech detection:
  • $\triangleright$
    Size: datasets should contain more than 5,000 samples to ensure adequate size for subsequent sampling.
  • $\triangleright$
    Label: datasets should contain diverse label space that address different facets of offensive language. Both neutral and non-neutral categories should be included for parity.
  • $\triangleright$
    Definition: each dataset should have the associated definitions with each label available. This criteria is important to generate informative explanations.
  • $\triangleright$
    Target / Rationale: the dataset should elicit the targets and/or reasoning to justify their choice of label (rationale) of offensive sentiments, whether through discrete labels or free text.
These criteria significantly reduce the number of eligible candidates since many do not contain information on target and the associated definitions. Table 0(a) lists the 8 selected datasets.
Datasets for Evaluation.
Using similar criteria, we also select 3 datasets with different label spaces and definitions to test downstream models’ generalizability (details in Table 0(b)).
HateCheck was created by Röttger et al. (2021) with the explicit goal of evaluating hate speech detection models. The authors recruited and trained 10 annotators on label the cases on a binary schema: Hateful or Non-hateful, and reported an almost perfect final inter-annotator agreement of 0.93.
HateXplain was constructed by Mathew et al. (2021). The authors asked annotators to highlight the span of tokens, called rationales, that contribute to their selection of the labels: Hate, Offensive, Normal. This dataset is primarily collected from Twitter and the Gab platform.
Latent_Hate was created by ElSherief et al. (2021a) on the premise that offensive speech classifiers tend to bias towards covert negative sentiment. With 22, 584 samples, this dataset is collected from Twitter( we discard the augmentation from the Social_Bias dataset avoid contaminating the training corpus (Sap et al., 2019)) . This set contains 2 coarse-grained labels Not Hate, Explicit Hate, Implicit Hate, while a subset contains 6 fine-grained categories, which we refer to as Implicit_Hate.
Dataset Total Size Sample Size Platform Target Ration. Selected Labels
Salminen et al. (2018)
137,098
5, 418
Y, F
Hateful, Neutral
Qian et al. (2019)
34,000
5, 034
G
Not Hate, Hate
Sap et al. (2019)
44,671
6, 033
G, R, T
Not Offensive, Offensive
Vidgen et al. (2021a)
27,494
6, 717
R
Neutral, Person Directed Abuse, Affiliation Directed Abuse, Identity Directed Abuse
Vidgen et al. (2021b)
10,152
7, 209
S
None, Derogation, Dehumanization, Animosity, Support, Threatening
Sachdeva et al. (2022)
135,556
7, 272
Y, T, R
Not Hate speech, Hate Speech
Hartvigsen et al. (2022)
274,186
7, 239
S
Benign, Toxic
Toraman et al. (2022)
100,000
7, 215
T
Normal, Offensive, Hate
Total
52, 137
(a) Datases used to create training corpus. Sample Size denotes the number of chosen samples from corresponding dataset included in the training corpus.
Dataset Total Size Split Ratio K val K test Platform Target Ration. Selected Labels
HateCheck
3,728
50:50
300
500
S
Non-hateful, Hateful
HateXplain
20,148
60:40
200
400
G, T
Normal, Offensive, Hate
Latent_Hate
19,112
60:40
200
400
T
Not Hate, Explicit Hate, Implicit Hate
Implicit_Hate
4,153
60:40
150
T
White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation
(b) Datasets for testing
Table 1: Sample Size denotes the number of entries in the final corpus. Target and Ration. indicates the availability of annotation on Target or Rationale in the dataset. For Platform, F: Facebook, Y: Youtube, G: Gab, R: Reddit, S: Synthetic, T: Twitter. K val and K test represent the number of sampler per class drawn during development of the training corpus and final testing, respectively. Full definitions in Table 6 and 7.
Obtaining Annotation-Guided Explanation.
Yang et al. (2023) showed that using GPT-3 to augment human-written rationales creates linguistically coherent while maintains faithfulness to the original content. We use the template shown in Figure 1 to generate the explanation for each input guided by the available annotations on label, target, and rationale from the chosen corpora. For datasets that contains multiple annotations for each sample, we select the ultimate label via majority voting and concatenate individual annotations on the targets and/or rationales into a single string separated by the delimiter "|".
Figure 1: Template used to obtain explanations from GPT-3.5-Turbo guided by human-annotated rationales.
We choose GPT-3.5-Turbo accessed via the OpenAI’s API to generate explanations due to the balance between the model’s affordability and its capability to follow instructions and generate coherent outputs (Ye et al., 2023a; Koubaa, 2023). For each of the 8 training datasets, we first randomly select and qualitatively analyze 20 samples to ensure the generated explanations are a) stylistically coherent, b) consistent with the provided labels, and c) congruent with the criteria denoted by the definitions. If deemed unsatisfactory, we iteratively adjust the input prompt until the quality threshold is achieved. Appendix A.3 reveals more specific details on the final prompts for different scenarios.
3.2 Optimization of Synthesized Corpus
Extending previous works Magister et al. (2022); Ho et al. (2022), we are interested in optimizing 2 parameters central to the construction of our corpus: the distribution of neutral vs. non-neutral classes in the data and the number of explanations per sample. The former has been noted to influence predictive powers Rathpisey and Adji (2019); Casula and Tonelli (2020). The latter, also referred to as degree of reasoning diversity, could improve knowledge distillation Ho et al. (2022).
3.2.1 Setup
Below are the experiments we perform on a sample of collected data to optimize these 2 parameters.
Description of Procedure.
We choose 1,000 samples from each of the eight training datasets based on the following distribution: 20% are selected from neutral samples (categories that do not indicate any offensive content, e.g., Not Hate, Normal), while the remaining samples are evenly distributed among the non-neutral categories. Inspired by Ho et al. (2022) that diverse reasoning paths could improve knowledge distillation, we collect 4 alternative explanations, or degree of reasoning diversity, generated by GPT-3.5-Turbo for these samples using temperature 0.7, resulting in 32,000 samples.
Figure 4 illustrates the Alpaca-styled template to format each post with its corresponding label, generated explanation, definitions along with the instruction into blocks of an input prompt (Taori et al., 2023). Using the described corpus, we supervised finetune the open-sourced model Llama 2 Chat-HF of 7 billion parameters (hereby referred to as Llama 7B) (Touvron et al., 2023) via LoRA techniques (Technical details specified in Appendix A.2) (Hu et al., 2021). Then, we perform zero-shot classification using the same template to prompt the fine-tuned models to generate the explanation and label for posts drawn from the 3 test datasets. For this part, posts are drawn using K-shot sampling (equal number of samples for each class) on the Validation set of the test data (K val in Table 1(b)). We omit Implicit_Hate due to this set’s markedly different 6-label space.
Experiment Configurations.
For the first experiment on the ratio of neutral vs. non-neutral categories, the training data is split into subsets with their respective distribution between the neutral (NE) class and non-neutral class(es) (NN) described by the following formula: $NN=R*NE$, where $R\in\{1,2,3,4\}$ is the ratio coefficient. We set the number of explanations per sample to 2 for this part. For the second experiment, we construct the subsets by varying the degree of reasoning diversity $D\in\{1,2,3,4\}$, of each post.
Answer Extraction.
We extract the explanation and predicted labels after their respective tag in the generated output. If the models generate multiple items from the dataset’s label space, we select the first legal label. If no acceptable output is obtained, we randomly select an item in the label space.
3.2.2 Insights and Augmentation
We report the model’s macro F1-scores on the validation set of each configuration in Table 2. A balanced distribution between neutral and non-neutral classes in the training corpus is beneficial, as reflected by the substantially high mean F1-score of 0.643 when R=1. On the other hand, having 3 explanations per sample (D=3) achieves the best overall performance across 3 test sets, consistent with Ho et al. (2022)’s findings on the benefit of multiple reasoning paths. Nevertheless, performance markedly degrades when D=4. Based on manual analysis, we hypothesize that the quality of generated outputs degrades as the degree of diversity increases, consequently affecting the performance of models trained on this data. These insights are valuable to practitioners when curating synthetic data from LLMs.
From these results, we augment the training corpus by selecting an additional approximately 1,800 samples from each of the 8 datasets while preserving the 1:1 balanced ratio of neutral to non-neutral classes. Then we collect 3 explanations per sample using the described mechanism, resulting in a final corpus of 52,137 samples (Table 0(a)), hereby referred to as the HateCOT corpus.
R=1 R=2 R=3 R=4
HateCheck 0.879 0.750 0.650 0.574
HateXplain 0.534 0.533 0.495 0.528
Latent_Hate 0.516 0.473 0.456 0.408
Average 0.643 0.585 0.534 0.503
(a) Results for Ratio configurations (R)
D=1 D=2 D=3 D=4
HateCheck 0.851 0.879 0.864 0.783
HateXplain 0.549 0.534 0.597 0.607
Latent_Hate 0.480 0.516 0.477 0.465
Average 0.627 0.643 0.646 0.618
(b) Results for number of explanations per sample (D)
Table 2: Macro F1-scores for different configurations of distribution between neutral vs. non-neutral classes (top) and number of explanations per sample (bottom) on validation set. Best average performance in bold.
4 Experiments on Test Sets
We perform experiments to answer 3 questions. First, does HateCOT improve zero-shot classification of open-sourced LLMs on unseen datasets? Second, how much data is necessary to enable competitive performance via in-domain fine-tuning after pretraining on HateCOT? Finally, is in-context learning a viable alternative to finetuning?
Models In addition to Llama 7B in Section 3.2.1, the following open-sourced models are selected.
  • $\triangleright$
    Llama 13B A larger variant of the instruction-tuned Llama 7B with 13 billion parameters.
  • $\triangleright$
    OPT-IML Based on the original OPT (Open Pre-trained Transformer Language Models) (Zhang et al., 2022), this encoder-only model contains 1.3 billion parameters and was further trained on the Instruction MetaLearning (IML) dataset Iyer et al. (2022).
  • $\triangleright$
    Flan-T5-L Chung et al. (2022) further instruction-finetuned the encoder-decoder T5 family of models (Raffel et al., 2020). We use the Large version of 780 million parameters.
  • $\triangleright$
    COT-T5-XL This model is a variant of the Flan-T5-XL (3 billion parameters) model that is further finetuned on the CoT dataset, a collection of 1.8 million samples augmented with chain-of-thought-style explanations (Kim et al., 2023).
4.1 Zero-shot Classification
This setting involves prompting the models to perform the classification task with no in-context examples via 2 modes: No Explanation, which prompts the model to directly predict the label for the given input, and With Explanation, which requires a justification before the predicted label. We finetune the base models using only HateCOT and evaluate their performance on the 3 test sets as in Section 3.2.2. Relevant technical details on pre-training are presented in Appendix A.2.
With results presented in Figure 1(a), we observe that the smaller Flan-T_L and OPT-IML are not able to generate explanations even when prompted to. In contrast, their scaled-up counterparts could follow instructions at all settings, indicating that size matters to effectively exploit instruction-tuning Wei et al. (2021)). Asking Base (off-the-shelve) models to provide explanations before label provides observable boost to Llama models in HateCheck and HateXPlain datasets, but actually hampers performance in Latent_Hate (and its derivative Implicit_Hate), a dataset noted for to be challenging for classifiers due to its covert nature ElSherief et al. (2021b).
Finetuning on HateCOT unanimously enables all models to generate explanations. While smaller models receive no observable boost in their performance, larger models (COT-T5-XL, Llama 7B, Llama 13B) are considerably enhanced compared to their base counterparts. For COT-T5-XL, with the exception of HateXplain, HateCOT-trained version attain F1 scores increment of 7.6% (HateXplain and 9.5% (Latent_Hate) over the base counterpart without explanations. Llama 7B observes 23.9%, 25.6%, 10% increment for HateCheck, HateXplain, Latent_Hate, respectively. These statistics are 27.9%, 118.5%, and 10.2% for Llama 13B. For Implicit_Hate, all models struggle to attain satisfactory performance, with COT-T5-XL leading the pack. Llama models’ performance improve with HateCOT, but still not much better than chance.
(a) Macro F-1 scores of LLMs in zero-shot setting using 3 configurations. Base refers to out-of-the-box models, +HateCOT denotes their finetuned counterpart on our dataset. FT5L: Flan-T5-L, CT5XL: COT-T5-XL. Results for Base Flan-T5-L and OPTIML models for With Explanation settings omitted to reflect their inability to generate explanation according to the prompt.
(b) Macro F-1 scores for models in zero-shot setting with explanation after K-shot in-domain finetuning at various values of K. Dashed line represents finetuned base models, solid line represents models pre-trained on HateCOT, then in-domain finetuned.
(c) Min, max and mean of macro F-1 scores for Llama 7B and Llama 13B over 5 seed using in-context learning. Pretrained denotes models finetuned on HateCOT only. Finetuned denotes ICL performed on models that have been both HateCOT-pretrained then K=256 shot in-domain finetuned.
Figure 2: Performance resutls of LLMs on test sets in various settings.
4.2 In-Domain Finetuning
We further finetune COT-T5-XL, Llama 7B, Llama 13B using data from the training set of the test datasets, including the Implicit_Hate set with 6-way classification. To simulate low-resource settings, we select 256 samples uniformly at random from each class of labels from the datasets, then augment them with explanations in the same fashion as described in Section 3.1. We perform K-shot sample (K $\in\{32,64,128,256\}$) to finetune and evaluate the models on the test sets. This process applies to both Base and Pretrained versions.
[…]
Interestingly, decoder-only models (Llama) considerably outperform encoder-decoder COT-T5-XL on the 2 and 3-way classification tasks, yet the reverse is observed for the nuanced 6-way Implicit_Hate. In fact, only COT-T5-XL consistently scales with the increment in training data to reach max F-1 score of 0.56, while Llama models plateau even at K=256 on this task at sub-0.3 range.
4.3 In-context Learning
As an alternative to in-domain fine-tuning, we investigate the models’ performance using in-context learning (ICL), when a number of complete examples are provided as part of the input prompt. We select 1 sample from each class in the training data of each dataset, then obtain its the associated explanation from GPT-3.5-Turbo. The sets of post, explanation and label are arranged in the same format in the same template as in Figure 4. We run inference for classification results over 5 random seeds, which also randomly permutes the samples’ order.
Figure 1(c) shows the mean, minimum and maximum values macro F-1 scores over the seeds for the base, HateCOT-tuned only (pre-trained), and in-domain finetuned (K=256) versions of Llama 7B and Llama 13B. COT-T5-XL regularly generates overly repetitive outputs, and thus omitted. The range of F-1 scores are large regardless of settings, an observation in line with the variance of in-context learning in literature Lu et al. (2022); Dong et al. (2022). Unsurprisingly, base models’ performances tend to be inferior to their tuned counterpart. Finetuned models display superior performance than the other 2 counterparts on average. Interestingly, the max F-1 scores of finetuned models in ICL are not appreciably better than those in the zero-shot counterparts (Figure 1(b)). In contrast, except for Llama 7B on HateXplain, the best scores of pretrained models approach those of the finetuned models–particularly for Llama 13B.
This finding suggests another advantage of pre-training on HateCOT then maximize performance via ICL without additional in-domain finetuning, an area that has attracted growing attention Min et al. (2021); Wang et al. (2023); Ye et al. (2023b). Nevertheless, there exists the trade-off: ICL examples with explanations extend significantly the context length, and ICL inferencing takes considerably more time compared to zero-shot, making the latter more resource-efficient overall.
4.4 Overall Assessment
Our experiments demonstrate that HateCOT confers various benefits relating to generalizing offensive speech detection. Priming LLMs using HateCOT via supervised training significantly enhances zero-shot predictive power. Competitive performance is possible to obtained using significantly less in-domain training data, a bottleneck factor on resources in the traditional pipeline. Our findings suggest suggest a cost-efficient alternative approach: practitioners could obtain human-derived annotation and rationales for K-shot samples, then augment them to acquire explanations using LLMs as outlined. This process may be iterated until desired performance is reached according to some guiding criteria, minimizing the waste of resources.
5 Quality of Explanations
We analyze the quality of the explanations generated by various versions of 3 models COT-T5-XL, Llama 7B, Llama 13B. Both recent literature Liu et al. (2023); Li et al. (2023) and our validation experiment (reported in QA 5.1 below) corroborate OpenAI’s GPT-4’s high agreement with human annotators. Thus, we leverage GPT-4’s judgement to compare our models’ outputs.
5.1 QA 1: Base vs. Pretrained
We randomly select 500 samples per model whose explanations from both the Base and HateCOT-pretrained versions agree on the predicted label from all combined test sets. Using the template in Figure 5, we ask GPT-4 to select its preferred explanation for the purpose of content moderation based on fluency, soundness and the alignment with the definition of the chosen label. We present the explanations from these 2 versions in a randomized order for the same post. From this pool, we recruit a human annotator to perform the same labelling task on 150 random samples. Human decisions match GPT-4’s 68% of the times, indicating considerable agreement (Appendix A.4) .
Converting GPT-4’s ratings to an ordinal scale (0 if Base model is preferred, 1 otherwise), we perform the Wilcoxon rank sum test with the null hypothesis $H_{0}$: there is no significant difference between the number of times Pretrained or Base model is preferred over the other. Table 3 shows the overwhelming preference for explanations of the Pretrained models. That $p$-value $\ll$ $\alpha=0.05$ rejects the null hypothesis and provides evidence of HateCOT-induced improvement on the quality of explanation, particularly for COT-T5-XL.
Model Base Pretrained p-value
COT-T5-XL
27   (5.4%)
473 (94.6%)
Llama 7B
109 (21.8%)
391 (78.2%)
Llama 13B
122 (24.4%)
378 (75.6%)
Table 3: Number of times outputs of Base vs Pretrained models are preferred by GPT-4 across 500 samples. *** indicates $p$-value $\ll$ 0.05 by the Wilcoxon rank test.
Figure 3: Heatmap for the average rating of explanations by finetuned Model (x-axis) and Dataset (y-axis) on 3 criteria from 1 (least) to 5 (very). Overall indicates average scores aggregated over all datasets. Triplets of scores italicized and in bold are those whose $p$-value < 0.05 by one-way ANOVA test that compare ratings of 3 models across the dataset on that row. Italicized-only scores indicate marginal significance ($p$-value $\approx$ 0.07).
5.2 QA 2: Inter-model Comparison
Inspired by Wang et al. (2023); Lin et al. (2023); Yang et al. (2023), we assess the quality of explanations generated by finetuned models (K=256) on the following criteria:
  • $\triangleright$
    Informativeness: whether the model provides factually correct and relevant information in its explanation.
  • $\triangleright$
    Persuasiveness: how convincingly the explanation justifies its chosen label for the post.
  • $\triangleright$
    Soundness: how valid and logical is the explanation with respect to the label’s definition.
Using the template in Figure 6, we collect numerical ratings from GPT-4 on a scale from 1 (least) to 5 (very) on these criteria for 200 posts per model per dataset, for a total of 2,400 samples. Figure 3 displays the mean ratings for each model-dataset pair, as well as Overall scores averaged across all datasets for each model. Ratings for Informativeness appear consistently lower than those of the other 2 criteria, likely due to the challenging context-dependent nature of generating post-specific relevant information. Overall, Persuasiveness and Soudness ratings around 4 out of 5 suggest high likelihood of acceptance by human users for content moderation.
We also note the general correlation between the models’ classification performance (Figure 1(b)) and mean ratings on each dataset. All finetuned models attain F-1 scores above 0.9 along with the highest average ratings on HateCheck, followed by HateXplain. Latent_Hate and its more difficult subset Implicit_Hate yield markedly lower average ratings. We perform ANOVA test on the ratings of models stratified by dataset and overall using $\alpha$=0.05. Aggregated ratings do not appear to differ meaningfully across criteria. Interestingly, COT-T5-XL’s superior performance over Llama models on Implicit_Hate coincides with statistically or marginally significant differences between average ratings. This correlation between better classification and higher quality of explanation is useful to calibrate models for real-life deployment.
5.3 QA 3: Target Identification
Dataset COT-T5-XL Llama 7B Llama 13B
HateCheck 94.5 % 96.5 % 93.5 %
HateXplain 71.0 % 74.0 % 74.0 %
Latent_Hate 50.5 % 51.0 % 44.5 %
Table 4: Percentage out of 200 samples per dataset, where explanations correctly identify at least 1 of the targets listed by human annotators by each model.
Finally, we investigate the in-domain finetuned models’ capabilities to identify the target of the sentiments expressed by the post. We also randomly select 200 samples from each dataset that have Target variable annotated by humans, then ask GPT-4 to judge whether the explanations from the models mention at least one of the listed targeted groups using the template in Figure 7. Note that we encourage GPT-4 to consider potential variance of expression and not restrict to exact matches.
Table 4 shows the percentage of accuracy on this task. Similar levels of descending performances are observed in the presented order of test datasets. However, this observation may be an artifact of the differences between the annotations targets among datasets: HateCheck has a limited number of discrete categories while Latent_hate contains multiple combinations of free-text labels. We urge practitioners to consider this factor while curating training data if target identification is desired.
6 Conclusion
We show empirically that our HateCOT dataset considerably enhances offensive speech detection even with limited training data while producing high-quality explanations. We invite researchers to build upon our work to address other related low-resource areas.
7 Limitations
We first acknowledge that our work is restricted to English corpora. Second, due to computational limitations, we could not perform experiments on larger open-source models. With the development of newer, more powerful models, it is reasonable to expect their performance to further improve though the use of this dataset of our corpus. Finally, an elaborated human study may further corroborate our quality analysis of explanations. Given supporting literature , we expect our findings to remain consistent Gilardi et al. (2023); Yang et al. (2023); Lin et al. (2023).
8 Ethics Statement
The authors note the risk of propagating implicit biases that LLMs are known to carry Cheng et al. (2023); Gupta et al. (2023). We do acknowledge the potential malicious usage of our corpus to generate content capable of evading detection, or jeopardizing classifiers’ performance.
Appendix A Appendix
A.1 Data Pre-Processing
The datasets used in this work are released by their respective authors for research purpose. Aware of the risk of containing confidential in social media data, we anonymize posts during the curation process by replacing user handles with the string ’<user>’. Our many layers of randomization provides further protection with respect to privacy.
A.2 Technical Specification
A.2.1 Inferencing
We used OpenAI’s API to access the publicly available version of GPT-3.5 in November 2023, and GPT-4 in January 2024. To obtain explanations from the former (as described in 3.2.2), we set the temperature to 0.7 during inference to obtain multiple outputs per sample. For GPT-4 and other open-source models, we use greedy decoding.
A.2.2 Finetuning
To train models, we employ both full supervise finetuning (FLAN-T5-Large, OPT-IML) and LoRA parameter-efficient techniques (all other models). LoRA models set to 8-bit quantization using the BitsandByes library. Training FLAN-T5-Large, OPT-IML, LLAMA 7B models was done on 2 Nvidia RTX A6000 GPUs, whereas COT-T5-XL, Llama 13B used 4 GPUs. Hyperparameters for the following candidates are tuned on the validation set of sampling K=64 shots from the leftover training samples by optimizing macro F1-score metric. Options in bold indicate final chosen values among multiple across all models to finetune on HateCOT.
  • Learning rate: {5e-5, 1e-4, 3e-4}
  • Training Epochs: {1, 2, 3}
  • LoRA Rank: {16, 32, 64} (alpha===rank*2)
  • LoRA Target Modules: {Q, V}
  • Batch size: 2
  • Gradient Accumulation Step: 2
For in-domain K-shot finetuning, the values above remain the same except for the following variations in Learning Rates, which is set to 1e-4 for HateCheck, HateXplain, Latent_Hate, and 3e-4 otherwise.
A.3 Variants of Prompt Template for Explanation
The 8 datasets introduced in Section 3.1 may not always have annotations on all required fields; thus, we modify the first sentence in the Instruction block in Figure 4 with the following variants when appropriate:
  • Only Target is available: ’Provide a brief paragraph to explain step-by-step how the post targets the specified group or entity, and how that leads to the specified Label based on the given Definitions.’
  • Both Target and Rationale are available: ’By elaborating on the provided Annotation, provide a brief paragraph to explain step-by-step how the post targets the specified group or entity, and how that leads to the specified Label based on the given Definitions.’
Model Base Pretrained
COT-T5-XL
11 (22%) 39 (78%)
Llama 7B
20 (40%) 30 (60%)
Llama 13B
23 (46%) 27 (54%)
Table 5: Number of times human annotator prefers outputs of Base vs Pretrained in Section 5.1
A.4 Human Annotation for QA Experiment
With approved IRB, we recruit a native English speaker from the United States using the Prolific platform to annotate 50 samples per model, for a total of 150 data points for the task described in Section 5.1. The annotator was paid a fair wage at \$15 per hour, and forewarned about the nature of the task. Table 5 shows the results. We observe that preference for explanations generated by Pretrained models remains consistent with GPT-4’s.
Dataset Definition
Salminen et al.,2018
Neutral : A post that is not offensive to any group of people. Hateful: An offensive post, motivated, in whole or part, by the writer’s bias against an aspect of a group of people.
Qian et al.,2019
Not Hate : Does not contain any attack on people based on protected characteristics. Hate: Contains direct attack on people based on protected characteristics such as race, ethnicity, national origin, religious affiliation, sexual orientation, caste, sex, gender, gender identity, and serious disease or disability.
Sap et al.,2019
Not Offensive : not offensive to anyone. Offensive : denotes the overall rudeness, disrespect, or toxicity of a post. whether a post could be considered offensive to anyone.
Vidgen et al.,2021a
Neutral: Content that does not fall into other categories, usually entirely unrelated to abuse, hate, prejudice, or intolerance. Identity Directed Abuse: Content that directs abuse at an identity, which relates to fundamental aspects of individuals’ social position, community and self-representation. An identity includes but is not limited to religion, race, ethnicity, gender, sexuality and sexual preferences, immigration status, nationality, ableness, physical appearance and class. Affiliation Directed Abuse : Content that directs abuse at people who have a voluntary affiliation with a profession, membership, association, ideology, or other well-defined group or collective. Person Directed Abuse : Content that directs abuse at an identifiable person.
Vidgen et al.,2021b
Derogation : content which explicitly attacks, demonizes, demeans or insults a group. Animosity : content which expresses abuse against a group in an implicit or subtle manner. Threatening : content which expresses intention to, support for, or encourages inflicting harm on a group, or identified members of the group. Support For Hateful Entities : content which explicitly glorifies, justifies or supports hateful actions, events, organizations, tropes and individuals collectively, entities. Dehumanization : content which perceives or treats people as less than human. Not Hate : content that falls into none of the other categories.
Basile et al.,2019
Hate Speech : language that is used to expresses hatred towards a targeted group or is intended to be derogatory, to humiliate,or to insult the members of the group. may also be language that threatens or incites violence. Offensive Language : may contain offensive terms but targets disadvantaged social groups in a manner that is potentially harmful to them. Neither : language that does not all into either of the other categories.
Hartvigsen et al.,2022
Benign: Text that is not harmful nor offensive to anyone. Toxic: Text that could be seen as harmful to anyone, may contain offensive, rude, humor, insults, personal attacks, profanity, aggression, may refer to targeted group with harmful intent that is expressed in stereotypes or lewd manners
Toraman et al.,2022
Hate : target, incite violence against, threaten, or call for physical damage for an individual or a group of people because of some identifying trait or characteristic. Offensive : humiliate, taunt, discriminate, or insult an individual or a group of people in any form, including textual. Normal : does not fall into any of the other categories.
Table 6: Labels and definitions of 8 datasets used to synthesize training corpus.
Dataset Definition
HateCheck
Non-hateful: Language that does not contain any abusive sentiment targeted at a protected group. Hateful: Abuse that is targeted at a protected group or at its members being a part of that group. Protected groups include age, disability, gender identity, familial status, pregnancy, race, national or ethnic origins, religion, sex or sexual orientation.
HateXplain
Normal : neither hate speech nor offensive. Hate Speech : language which attacks, demeans, offends, threatens, or insults a group based on race, ethnic origin, religion, disability, gender, age, sexual orientation, or other traits. it is not the presence of certain words that makes the text hate speech, rather you should look the context the word is used in the text. Offensive Language : usage of rude, hurtful, derogatory, obscene or insulting language to upset or embarrass people.
Latent_Hate
Not Hate : speech or actions that do not involve any form of hatred, prejudice, or discrimination toward individuals or groups based on their characteristics. Explicit Hate: openly expressed, direct forms of hatred and prejudice toward individuals or groups based on their characteristics. Implicit Hate: coded or indirect language that disparages a person or group on the basis of protected characteristics like race, gender, and cultural identity.
Implicit_hate
White Grievance : includes frustration over a minority group s perceived privilege and casting majority groups as the real victims of racism. this language is linked to extremist behavior and support for violence. Incitement To Violence : includes flaunting in group unity and power or elevating known hate groups and ideologies. Inferiority Language : implies one group or individual is inferior to another, and it can include dehumanization denial of a person s humanity , and toxic language that compares the target with disease, insects, animals . related to assaults on human dignity, dominance, and declarations of superiority of the in group. Irony : refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. Stereotypes And Misinformation : associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like holocaust denial and other forms of historical negationism. Threatening And Intimidation : conveys a speaker’s commitment to a target s pain, injury, damage, loss, or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation.
Table 7: Labels and definitions for 4 test datasets.
Figure 4: Template used to prompt LLM for classification inference.
Figure 5: Template for QA Experiment 1. In this example, EXP_A is preferred.
Figure 6: Template for QA Experiment 2.
Figure 7: Template for QA Experiment 3.