### v4

# Instructions for *ACL Proceedings

Abstract
Warning: This paper contains examples of very offensive material. The widespread use of social media necessitates reliable and efficient detection of offensive content to mitigate harmful effects. Although sophisticated models perform well on individual datasets, they often fail to generalize due to varying definitions and labeling of "offensive content." In this paper, we introduce HateCOT, an English dataset with over 52,000 samples from diverse sources, featuring explanations generated by GPT-3.5-Turbo and curated by humans. We demonstrate that pretraining on HateCOT significantly enhances the performance of open-source Large Language Models on three benchmark datasets for offensive content detection in both zero-shot and few-shot settings, despite differences in domain and task. Additionally, HateCOT facilitates effective K-shot fine-tuning of LLMs with limited data and improves the quality of their explanations, as confirmed by our human evaluation. Our repository is available at https://github.com/hnghiem-usc/hatecot .
HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models
Huy Nghiem University of Maryland nghiemh@umd.edu                        Hal Daumé III University of Maryland & Microsoft Research hal3@umd.edu
1 Introduction
As social media has become indispensable to modern discourse, this channel of communication has amplified the propagation of offensive content. Speech that promotes hateful sentiments thrives on such platforms, leading to real and significant harm on their audience Giachanou and Rosso (2020); Saha et al. (2019). However, “offensive content” is still a contested construct, as what is and is not allowed varies by platform. In research, different approaches analyze semantically similar but still distinct concepts: Cyber-bullying, Toxicity, Sexist, Racist, Hate etc. Poletto et al. (2021); Fortuna et al. (2021); Nghiem et al. (2024); Nguyen et al. (2023), further highlighting this contestedness.
Compounding the challenge, reliable detection of offensive content typically requires significant amounts of data. Sophisticated models tend to be data-hungry, and the process of curating a dataset tailored to a specific use case can be costly, time-consuming, and emotionally challenging for annotators Founta et al. (2018); Toraman et al. (2022). The typical pipeline consists of collecting samples based on topic-relevant key words, then recruiting either crowdworkers or experts to annotate data before developing classification models Paullada et al. (2021). Each step incurs investment and may inject subtle idiosyncrasies proportionate to the size of the downstream dataset, further limiting transferable usefulness to related tasks Fortuna et al. (2021). The size of a dataset also does not necessarily guarantee cross-domain transferrability Poletto et al. (2021); Fortuna et al. (2021).
In practical settings, users may desire transparency from social media platforms. Therefore, the ability to provide human-understandable justification based on platform-specific policy becomes an attractive feature for content moderation. Nevertheless, current techniques often still fail to offer intuitive explanatory signals Yadav et al. (2023); Babaeianjelodar et al. (2022); Ibrahim et al. (2022).
In this work, we attempt to simultaneously reduce the cost of data curation, enhance cross-dataset generalization, and address the necessity of explainable decisions for offensive content detection. Specifically, our main contributions are:
  1. 1.
    We release HateCOT (Hate-related Chains-of-Thought), a dataset of over $52,000$ samples consisting of input text, a hate speech label, and an explanation for that label. This corpus is constructed by merging eight datasets with explanations created by using GPT-3.5-Turbo to augment human annotations.
  2. 2.
    We demonstrate the benefits of using HateCOT as a pretraining corpus before finetuning on a target domain. Empirical results across 3 datasets show that open-source Large Language Models (LLMs) can effectively leverage definitions to adapt to new tasks using zero-shot and few-shot settings via finetuning and in-context learning .
  3. 3.
    We assess the quality of explanations generated by our finetuned models with respect to the criteria described in their corresponding definitions. These insights showcase LLM-generated explanations as a means to enhance transparency in content moderation.
2 Related Works
2.1 Offensive Speech Detection
Offensive speech detection has attracted considerable interest from the research community. Earlier approach typically investigated coarse-grained labels (e.g. Hate vs Not Hate) while subsequent efforts explored more diverse facets of offensive speech at higher granularity Founta and Specia (2021); Poletto et al. (2021); Vidgen and Derczynski (2020). Increasingly more advanced models emerged over time with the diversity of datasets.
Cross-domain generalization, however, still remains a relevant challenge in the area. Fortuna et al. (2021) found empirically that cross-dataset transference is highly dependent on semantic similarity between their label spaces. Recent works have pretrained Transformer-based models, such as HateBERT and fBERT, on specialized corpora to enhance generalization to various levels of success Caselli et al. (2021); Sarkar et al. (2021).
2.2 LLMs in Offensive Speech Classification
Zampieri et al. (2023) assessed a range of open-sourced LLMs on zero-shot prompting on the OffensEval task and found their performance trailing by a wide margin behind trained existing finetuned BERT-based systems. Chiu et al. (2021) and Han and Tang (2022) used the proprietary GPT-3 on a set of different datasets and noted that informative contexts and examples could boost the model’s performance to competitive levels on a different set of data. Similarly, Roy et al. (2023) found that adding explanation to pipeline could result in 10 to 20% boost in performance of LLMs over baselines. (Yang et al., 2023)’s study found that training LLMs with step-by-step reasoning grounded by annotations could improve predictive power.
Pretrained language models have exhibited remarkable ability in text generation Celikyilmaz et al. (2020). Recent large-size LLMs such as GPT-3 and later models are capable of astoundingly fluent, convincing and knowledge-infused outputs Zhang et al. (2023). LLMs with hundreds of billions of parameters even exhibit reasoning capabilities Wei et al. (2022, 2021), leading to a flurry of research on prompting techniques to harness their prowess, such as Chain-of-thought (COT), Tree-of-thought etc. (Yao et al., 2023; Diao et al., 2023). An interesting line of research leverages LLMs to efficiently generate high volumes of synthetic data for tasks with training resource is scarce Puri et al. (2020); Bao et al. (2023); Whitehouse et al. (2023). We build upon these works to construct a dataset that can induce smaller LLMs to efficiently adapt to new categories of offensive content by leveraging their provided definitions.
3 Building HateCOT
We first describe the process to identify the candidate datasets from literature, and the procedure to obtain annotation-guided explanations from these samples (Section 3.1). We then perform a set of validation experiments to optimize the data’s parameters for downstream tasks before augmenting our corpus to its eventual size (Section 3.2).
3.1 Data Selection
Datasets for Training.
We use the following criteria to filter existing corpora related to offensive speech detection:
  • $\triangleright$
    Size: datasets should contain more than 5,000 samples to ensure adequate size for subsequent sampling.
  • $\triangleright$
    Label: datasets should contain diverse label space that address different facets of offensive language. Both neutral and non-neutral categories should be included for parity.
  • $\triangleright$
    Definition: each dataset should have the associated definitions with each label available (Figure 6). This criteria is important to generate informative explanations.
  • $\triangleright$
    Target / Rationale: the dataset should provide the a) targets and/or b) rationales, which are fragments of free texts by human annotators to convey some understanding of the corresponding post. Figure 4 shows several examples of rationales, demonstrating their unsuitability to be used as explanations in their native format.
These criteria significantly reduce the number of eligible candidates since many do not provide the required annotation on target and definition. Table 1(a) lists the 8 selected datasets.
Datasets for Evaluation.
Using similar criteria, we select 3 additional datasets with different label spaces and definitions for downstream testing (details in Table 1(b)).
HateCheck was created by Röttger et al. (2021) with the explicit goal of evaluating hate speech detection models. 10 trained annotators labelled the dataset using a binary schema: Hateful and Non-hateful, with the reported inter-annotator agreement coefficient to be 0.93.
HateXplain was primarily collected from Twitter and the Gab platform Mathew et al. (2021). In addition to the labels Hate, Offensive, Normal, annotators also provide justification for their selection by highlighting the span of tokens, called rationales, that contribute to their decision.
Latent_Hate was created on the premise that offensive speech classifiers tend to bias towards covert negative sentiment ElSherief et al. (2021). After discarding the augmented portion from the Social_Bias dataset to avoid contamination, 22,584 samples collected from Twitter remained. This dataset contains 3 coarse-grained labels Not Hate, Explicit Hate, Implicit Hate, while a subset contains 6 fine-grained categories, which we refer to as Implicit_Hate in subsequent test regimens.
Dataset Total Size Sample Size Platform Target Ration. Selected Labels
Salminen et al. (2018) 137,098 5, 418 Y, F Hateful, Neutral
Qian et al. (2019) 34,000 5, 034 G Not Hate, Hate
Sap et al. (2020) 44,671 6, 033 G, R, T Not Offensive, Offensive
Vidgen et al. (2021a) 27,494 6, 717 R Neutral, Person Directed Abuse, Affiliation Directed Abuse, Identity Directed Abuse
Vidgen et al. (2021b) 10,152 7, 209 S None, Derogation, Dehumanization, Animosity, Support, Threatening
Sachdeva et al. (2022) 135,556 7, 272 Y, T, R Not Hate speech, Hate Speech
Hartvigsen et al. (2022) 274,186 7, 239 S Benign, Toxic
Toraman et al. (2022) 100,000 7, 215 T Normal, Offensive, Hate
Total - 52, 137
(a) Datasets used to create training corpus. Sample Size denotes the number of chosen samples from corresponding dataset included in the training corpus.
Dataset Total Size Split Ratio K val K test Platform Target Ration. Selected Labels
HateCheck 3,728 50:50 300 500 S Non-hateful, Hateful
HateXplain 20,148 60:40 200 400 G, T Normal, Offensive, Hate
Latent_Hate 19,112 60:40 200 400 T Not Hate, Explicit Hate, Implicit Hate
Implicit_Hate 4,153 60:40 - 150 T White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation
(b) Datasets for testing
Table 1: Sample Size denotes the number of entries in the final corpus. Target and Ration. indicates the availability of annotation on Target or Rationale in the dataset. For Platform, F: Facebook, Y: Youtube, G: Gab, R: Reddit, S: Synthetic, T: Twitter. K val and K test represent the number of sampler per class drawn during development of the training corpus and final testing, respectively. Full definitions in Table 6 and 7.
Obtaining Annotation-Guided Explanation.
Inspired by Yang et al. (2023)’s work that shows that GPT-3.5 could augment human-written rationales to create coherent texts that are still faithful to the original content, we use the prompt template in Figure 1 to generate the explanation, which is guided by the available annotations on label, target, and rationale from the chosen corpora. For datasets that contain multiple annotations per sample, we select the ultimate label via majority voting and concatenate annotations on the targets and/or rationales into a single string delimited by "|".
Figure 1: Template used to obtain explanations from GPT-3.5-Turbo guided by human-annotated rationales.
We use GPT-3.5-Turbo, accessed via the OpenAI’s API, to generate explanations due to this model’s affordability and its capability to follow instructions and generate coherent outputs (Ye et al., 2023a; Koubaa, 2023). For each of the 8 training datasets, we first randomly select and qualitatively analyze 20 samples to ensure the generated explanations are a) stylistically coherent, b) consistent with the provided labels, and c) congruent with the criteria denoted by the definitions. If deemed unsatisfactory, we iteratively adjust the input prompt until the quality threshold is achieved. Appendix A.4 describes this quality assurance process and the final prompts for different scenarios.
3.2 Optimization of Synthesized Corpus
Extending previous works Magister et al. (2022); Ho et al. (2023), we are interested in optimizing 2 parameters central to the construction of our corpus: the distribution of neutral vs. non-neutral classes in the data and the number of explanations per sample. The former has been noted to influence predictive powers Rathpisey and Adji (2019); Casula and Tonelli (2020). The latter, also referred to as degree of reasoning diversity, could improve knowledge distillation Ho et al. (2023). We use the open-source model Llama 2 Chat-HF of 7 billion parameters (hereby referred to as Llama 7B) (Touvron et al., 2023) to perform tuning experiments in this stage due to its manageable size and strong classification performance. These empirical findings then guide the final augmentation process.
3.2.1 Optimization Procedure
Below are the experiments we perform on a sample of the collected data to optimize these parameters.
Description of Procedure.
We choose 1,000 samples from each of the eight training datasets based on the following distribution: 20% are selected from neutral samples (categories that do not indicate any offensive content, e.g., Not Hate, Normal), while the remaining samples are evenly distributed among the non-neutral categories. Inspired by Ho et al. (2023) that diverse reasoning paths could improve knowledge distillation, we collect 4 alternative explanations, or degree of reasoning diversity, generated by GPT-3.5-Turbo for these samples using temperature 0.7, resulting in 32,000 samples.
Figure 5 illustrates the Alpaca-styled template to format each post with its corresponding label, generated explanation, definitions along with the instruction into blocks of an input prompt (Taori et al., 2023). Using the described corpus, we supervised finetune Llama 7B via LoRA techniques (technical details specified in Appendix A.2) (Hu et al., 2021). Then, we perform zero-shot classification using the same template to prompt the finetuned model to generate the explanation and label for posts drawn from the test datasets HateCheck, HateXplain, Latent_Hate. We omit Implicit_Hate at this stage due to this set’s markedly different 6-label space. For this part, posts are drawn using K-shot sampling (equal number of samples for each class) on the Validation portion of the test data, based on the values of K val shown in Table 2(b).
Experiment Configurations.
For the first experiment, the training data is split into subsets whose distribution between the neutral (NE) class and non-neutral class(es) (NN) described by the following formula: $NN=R*NE$, where $R\in\{1,2,3,4\}$ is the ratio coefficient. We set the number of explanations per sample to 2 , the smallest value that still enables the benefit of reasoning diversity. For the second experiment, we construct the subsets by varying the degree of reasoning diversity $D\in\{1,2,3,4\}$ of each post.
Answer Extraction.
We extract the generated explanation and predicted labels after their respective tags. If the models generate multiple items from the dataset’s label space, we select the first admissible label. If no acceptable output is obtained, we randomly select an item in the label space.
3.2.2 Insights and Augmentation
We report Llama 7B’s macro F1-scores on the validation set of each configuration in Table 2. A balanced distribution between neutral and non-neutral classes in the training corpus is beneficial, as reflected by the substantially high mean F1-score of 0.643 when R=1. On the other hand, having 3 explanations per sample (D=3) achieves the best overall performance across 3 test sets, consistent with Ho et al. (2023)’s findings on the benefit of multiple reasoning paths. However, performance markedly degrades when D=4. Our manual analysis reveals that the quality of generated outputs deteriorates as the degree of diversity increases, consequently affecting the performance of models trained on this data.
Guided by these empirical findings, we augment the training corpus by selecting approximately 1,800 extra samples from each of the 8 datasets while preserving the 1:1 balanced ratio of neutral to non-neutral classes. Then, we collect 3 explanations per sample using the described mechanism, resulting in a final corpus of 52,137 samples (Table 1(a)), hereby referred to as HateCOT.
R=1 R=2 R=3 R=4
HateCheck 0.879 0.750 0.650 0.574
HateXplain 0.534 0.533 0.495 0.528
Latent_Hate 0.516 0.473 0.456 0.408
Average 0.643 0.585 0.534 0.503
(a) Results for Ratio configurations (R)
D=1 D=2 D=3 D=4
HateCheck 0.851 0.879 0.864 0.783
HateXplain 0.549 0.534 0.597 0.607
Latent_Hate 0.480 0.516 0.477 0.465
Average 0.627 0.643 0.646 0.618
(b) Results for number of explanations per sample (D)
Table 2: Macro F1-scores for different configurations of distribution between neutral vs. non-neutral classes (top) and number of explanations per sample (bottom) on validation set. Best average performance in bold.
4 Experiments on Test Sets
We perform experiments to answer 3 questions. First, does HateCOT improve zero-shot classification of open-sourced LLMs on unseen datasets? Second, how much data is necessary to enable competitive performance via in-domain finetuning after pretraining on HateCOT? Finally, is in-context learning a viable alternative to finetuning?
Models In addition to Llama 7B in Section 3.2.1, the following open-sourced models are selected.
  • $\triangleright$
    Llama 13B A larger variant of the instruction-tuned Llama 7B with 13 billion parameters.
  • $\triangleright$
    OPT-IML Based on the original OPT (Open Pre-trained Transformer Language Models) (Zhang et al., 2022), this encoder-only model contains 1.3 billion parameters and was further trained on the Instruction MetaLearning (IML) dataset Iyer et al. (2022).
  • $\triangleright$
    Flan-T5-L Chung et al. (2024) further instruction-finetuned the encoder-decoder T5 family of models (Raffel et al., 2020). We use the Large version of 780 million parameters.
  • $\triangleright$
    COT-T5-XL A variant of the Flan-T5-XL (3 billion parameters), this model is further finetuned on the CoT dataset, a collection of 1.8 million samples augmented with chain-of-thought-style explanations (Kim et al., 2023).
4.1 Zero-shot Classification
We prompt the models to perform classification with no in-context examples via 2 modes: No Explanation, where the model directly predicts the label for the input, and With Explanation, where a justification is required before the predicted label. We finetune the base models using only HateCOT and evaluate their performance on the 4 test sets as in Section 3.2.2 (more details in Appendix A.2).
From results presented in Figure 2(a), the smaller models Flan-T5-L and OPT-IML are unable to generate explanations when prompted. In contrast, their scaled-up counterparts could follow instructions at all settings. Asking base (off-the-shelve) models to generate an explanation before the label results in observable boost to Llama models on HateCheck and HateXPlain, but actually hampers performance on Latent_Hate and its derivative Implicit_Hate, which are notably challenging due to its covert nature ElSherief et al. (2021).
Model Choice Matters
Pretraining on HateCOT unanimously enables all models to generate explanations. While smaller models receive no observable boost, larger models (COT-T5-XL, Llama 7B, Llama 13B) are considerably enhanced compared to their base counterparts. With the exception of HateXplain, the HateCOT-pretrained version of COT-T5-XL attains an increment in F1 scores of 7.6% on HateXplain and 9.5% on Latent_Hate over the base counterpart without explanations. Similarly, Llama 7B observes 23.9%, 25.6%, 10% increment on HateCheck, HateXplain and Latent_Hate, respectively. These statistics are 27.9%, 118.5%, and 10.2% for Llama 13B. Notably, all models yield non-competitive performance on Implicit_Hate.
The reduced performance of HateCOT-pretrained models compared to their base counterparts without explanations (e.g.: Flan-T5-L and OPT-IML on almost all test sets) is in line with literature as COT-style prompting tends to favor larger models Wei et al. (2022); Wang et al. (2024); Suzgun et al. (2023). Even the reduced F1 scores of COT-T5-XL on HateXPlain and Implicit_Hate is consistent with this model’s suboptimal performance relative to its larger variants, as showcased in Kim et al. (2023). These results serve as an empirical reference for researchers to select the appropriate model size for their respective task.
(a) Macro F1 scores of LLMs in zero-shot setting using 3 configurations. Base refers to out-of-the-box models, +HateCOT denotes their pretrained counterpart on our dataset. FT5L: Flan-T5-L, CT5XL: COT-T5-XL. Results for Base Flan-T5-L and OPTIML models for With Explanation settings omitted to reflect their inability to generate explanation according to the prompt.
(b) Macro F1 scores for models in zero-shot setting with explanation after K-shot in-domain finetuning at various values of K. Dashed line represents finetuned base models, solid line represents models pre-trained on HateCOT, then in-domain finetuned. For each dataset, the horizontal dashed line represents the base version of the best performing model at K=256 which is finetuned using the entire training data without any rationale for comparison, denoted as Best model (Full data, No. exp).
(c) Min, max and mean of macro F1 scores for Llama 7B and Llama 13B over 5 seed using in-context learning. Pretrained denotes models finetuned on HateCOT only. Finetuned denotes ICL performed on models that have been both HateCOT-pretrained then K=256 shot in-domain finetuned.
Figure 2: Performance resutls of LLMs on test sets in various settings.
4.2 In-Domain Finetuning
We further finetune COT-T5-XL, Llama 7B, Llama 13B using data from the training portions of the test datasets, including Implicit_Hate. To simulate low-resource settings, we choose 256 samples uniformly at random from each class, then augment them with explanations as described in Section 3.1. Both the Base and Pretrained versions of the models are then finetuned using various K-shot $\in\{32,64,128,256\}$ training data from this pool.
[…]
Interestingly, decoder-only models (Llama) considerably outperform encoder-decoder COT-T5-XL on the 2 and 3-way classification tasks, yet the reverse is observed for the nuanced 6-way Implicit_Hate. On this task, only COT-T5-XL consistently scales with the increment in training data to reach the max F1 score of 0.56, while Llama models plateau at sub-0.3 range even at K=256.
We further select the best performing model at K=256 for each dataset and finetune their Base versions with the entire training data and no explanations no definition. In Figure 2(b) and Table 3, in-domain finetuning after HateCOT achieves competitive results even with only a fraction of the full training data. Furthermore, prompting for explanations enables Llama 13B and COT-T5-XL to attain performance that surpasses using the full training data on Latent_Hate and Implicit_Hate.
4.3 In-context Learning
As an alternative to in-domain finetuning, we investigate the models’ performance using in-context learning (ICL), when a number of complete examples are provided as part of the input prompt. We select 1 sample from each class in the training data of each dataset, then obtain its the associated explanation from GPT-3.5-Turbo. The sets of post, explanation and label are arranged in the same format in the same template shown in Figure 5. We run inference for classification results over 5 seeds, which also randomly permutes their order.
Figure 2(c) shows the mean, minimum and maximum values of macro F1 scores over the seeds for the base, HateCOT-pretrained only (Pretrained), and in-domain finetuned (K=256) versions of Llama 7B and Llama 13B. COT-T5-XL regularly generates overly repetitive outputs, and thus omitted. The range of F1 scores is large regardless of settings, an observation in line with the variance of in-context learning in literature Lu et al. (2022); Dong et al. (2022). Unsurprisingly, base models’ performances tend to be inferior to their finetuned counterpart. Interestingly, the max F1 scores of finetuned models with ICL are not appreciably better than those in the zero-shot counterparts (Figure 2(b)). In contrast, except for Llama 7B on HateXplain, the best scores of pretrained models approach those of the finetuned models–particularly for Llama 13B.
This finding suggests another advantage of pre-training on HateCOT: boosting performance via ICL without additional in-domain finetuning, an area that has attracted growing attention Min et al. (2021); Wang et al. (2023b); Ye et al. (2023b). Nevertheless, there exists the trade-off: ICL examples with explanations extend significantly the context length, and ICL inferencing takes considerably more time compared to zero-shot, making the latter more resource-efficient overall.
Dataset Best Model @K=256 F1 @K=256 F1 Base + Full F1 % Data Size @K=256 Data Size Full Data %
HateCheck LLAMA 13B 0.95 0.99 96% 512 1,864 27%
HateXplain LLAMA 13B 0.64 0.72 89% 768 12,088 6%
Latent_Hate LLAMA 13B 0.66 0.64 103% 768 11,460 7%
Implicit_Hate COT-T5-XL 0.56 0.38 147% 1,536 2,707 57%
Table 3: Comparison of performance metrics for the best performing models finetuned using K=256 in-domain post HateCOT vs. finetuned on the full training set and no explanation nor definition. F1% denotes the percentage of macro F1 score of the K=256 finetuned model over that of the model trained on full data. Similarly, Data % denotes the percentage of data size used by the K=256 regimen over the full data.
4.4 Assessment and Recommendations
From empirical observations, we make the following recommendations to construct a cost-efficient pipeline for classifier on novel domains:
  • $\triangleright$
    The most consistent benefit of HateCOT is its capacity to enable data-efficient in-domain finetuning following pretraining.
  • $\triangleright$
    Practitioners should choose models of sufficient number of parameters for the task. Larger instruction-tuned LLMs appear to more effectively capitalize on HateCOT pretraining regimen before in-domain finetuning.
  • $\triangleright$
    Instead of devoting resources to curate substantial training data, practitioners could focus on obtaining high quality annotations for representative rationales, and augment them into explanations using their LLM of choice. Alternatively, practitioners may choose to curate the explanations organically to achieve certain desired thematic qualities. This process may be iterated until targeted performance is reached according to some guiding metrics with acceptable quality of explanation.
5 Quality of Explanations
In addition to the enhanced classification prowess, we investigate whether pretraining on HateCOT also improves the quality of explanation LLMs. To this end, the following 2 human quality assurance (QA) experiments are conducted. In QA 1, we assess if human annotators prefer the explanations generated by the base or HateCOT-pretrained LLMs (COT-T5-XL, Llama 7B, Llama 13B). In QA 2, we perform in-domain K-shot finetuning on the aforementioned models and examine how annotators evaluate their generated explanations. An additional assessment of Target Identification is presented in Appendix A.6.
5.1 QA 1: Base vs. Pretrained
From the outputs of the test sets generated by the 3 LLMs, we select 50 samples uniformly at random whose explanations from the Base and HateCOT-pretrained versions agree on the predicted label, for a total of 150 samples and 300 explanations. We then recruit 13 annotators from the crowdsource platform Amazon Mechanical Turk and solicit their annotation on these explanations (Appendix A.5). Using the template in Figure 6, we ask the annotators to indicate their preferred explanation that better suits the purpose of content moderation based on fluency, soundness and the alignment with the definition of the chosen label. Each post is annotated by 5 humans, resulting in 750 annotations.
In Table 4, we observe that the raw frequency count for explanations generated by the HateCOT-pretrained models exceed their base version’s. Similarly, even when tallying by majority vote–where the explanation is chosen by at least 3 out of 5 annotators–preference for those generated by the Pretrained models still prevails. We note that the preference margin is smaller for Llama 13B Pretrained, likely due to this model’s already strong generative capabilities.
Human (Frequency Count) Human (Majority Vote)
Model Base Pretrained Base Pretrained
COT-T5-XL 62 (24.8%) 188 (75.2%) 6 (12%) 44 (88%)
Llama 7B 109 (43.6%) 141 (56.4%) 19 (38%) 31 (62%)
Llama 13B 114 (45.6%) 136 (54.4%) 22 (44%) 28 (56%)
Table 4: Comparison of Base and Pretrained models in Human Evaluation. Frequency Count : count per annotation; Majority Vote indicates aggregate count by the version is preferred by at least 3 out of 5 annotators.
5.2 QA 2: Inter-model Comparison
Inspired by Wang et al. (2023a); Lin et al. (2023); Yang et al. (2023), we assess the quality of explanations generated by finetuned models (K=256) on the following criteria:
  • $\triangleright$
    Persuasiveness: how convincingly the explanation justifies its chosen label for the post.
  • $\triangleright$
    Soundness: how valid and logical is the explanation with respect to the label’s definition.
Figure 3: Heatmap for the average rating of explanations by finetuned Model (x-axis) and Dataset (y-axis) on 3 criteria from 1 (least) to 5 (very). Overall indicates average scores aggregated over all datasets. Triplets of scores italicized and in bold are those whose $p$-value < 0.05 by one-way ANOVA test that compare ratings of 3 models across the dataset on that row. Italicized-only scores indicate marginal significance ($p$-value $\approx$ 0.07).
For this QA task, we recruit 6 annotators also from Amazon Mechanical Turk (Appendix A.5). Using the template in Figure 7, we collect their numerical ratings on a scale from 1 (least) to 5 (very) on these criteria for 50 posts per model per dataset, for a total of 600 annotations. Figure 3 displays the mean ratings for each model-dataset pair, as well as Overall scores averaged across all datasets. The average Ovearall ratings for both Persuasiveness and Soudness are above 3.2 out of 5, indicating generally positive reception by human evaluators. Interestingly, there exists a degree of correlation between the models’ better classification performance (Figure 2(b)) and higher mean ratings on each dataset, a useful artifact to calibrate models. These ratings may serve as benchmarks for future works.
6 Conclusion
We show empirically that our HateCOT dataset considerably enhances offensive speech detection even with limited training data while producing high-quality explanations. We invite future research to explore other benefits of LLM-augmented data and extend them to other related low-resource areas.
Limitations
We first acknowledge that our work is restricted to English corpora, a common limitation among literature on offensive speech Yin and Zubiaga (2021); Poletto et al. (2021). However, our approach sets a proof-of-concept for researchers to construct similar corpus in other languages by leveraging existing resources. Furthermore, our developmental pipeline is considerably more data-efficient than conventional approaches (Section 4.1, 4.2), potentially lowering the barrier of entry for practitioners without access to abundant resources. Therefore, this work invites further expansion on multilingual datasets, particularly to develop corpora with clearly defined definitions to facilitate synergy with other research.
Second, due to computational limitations, we could not perform experiments on larger open-source models. With the development of newer, more powerful models, it is reasonable to expect their performance to further improve though the use of this dataset of our corpus, as demonstrated by our empirical results.
Finally, we recognize the risk of propagating implicit biases that LLMs are known to carry Cheng et al. (2023); Gupta et al. (2023). However, we note that the approach of using LLMs (GPT-3.5-Turbo in this paper) to bridge the logical gaps in original rationales has been shown to produce outputs less prone to logical failures Yang et al. (2023). Biases in Pretrained Language Models have been attracting much attention in the research community. We invite further works to consider our approach to reduce hallucinations and biases in text generation.
Ethics Statement
We acknowledge the potential malicious usage of our corpus to generate content capable of evading detection, or jeopardizing classifiers’ performance.
Acknowledgement
This work is funded by the University of Maryland’s Institute for Trustworthy AI in Law & Society (TRAILS). We thank the service of ACL ARR reviewers, area chairs and the editors of the EMNLP conference for our paper’s publication.
Appendix A Appendix
A.1 Data Pre-Processing
The datasets used in this work are released by their respective authors for research purpose. Aware of the risk of containing confidential in social media data, we anonymize posts during the curation process by replacing user handles with the string ’<user>’. Our many layers of randomization provides further protection with respect to privacy.
A.2 Technical Specification
A.2.1 Inferencing
We used OpenAI’s API to access the publicly available version of GPT-3.5 in November 2023, and GPT-4 in January 2024. To obtain explanations from the former (as described in 3.2.2), we choose temperature among candidates $\{0.3,0.5,\textbf{0.7}\}$ and settle on the last value during inference. This value is selected based on literature and multiple iterations of qualitative analysis of outputs Yang et al. (2023); Kim et al. (2023). For GPT-4 and other open-source models, we use greedy decoding.
A.2.2 Finetuning
To train models, we employ both full supervise finetuning (FLAN-T5-Large, OPT-IML) and LoRA parameter-efficient techniques (all other models). LoRA models set to 8-bit quantization using the BitsandByes library. Training FLAN-T5-Large, OPT-IML, LLAMA 7B models was done on 2 Nvidia RTX A6000 GPUs, whereas COT-T5-XL, Llama 13B used 4 GPUs. Hyperparameters for the following candidates are tuned on the validation set of sampling K=64 shots from the leftover training samples by optimizing macro F1-score metric. Options in bold indicate final chosen values among multiple across all models to finetune on HateCOT.
  • Learning rate: {5e-5, 1e-4, 3e-4}
  • Training Epochs: {1, 2, 3}
  • LoRA Rank: {16, 32, 64} (alpha===rank*2)
  • LoRA Target Modules: {Q, V}
  • Batch size: 2
  • Gradient Accumulation Step: 2
For in-domain K-shot finetuning, the values above remain the same except for the following variations in Learning Rates, which is set to 1e-4 for HateCheck, HateXplain, Latent_Hate, and 3e-4 otherwise.
A.3 Variants of Prompt Template for Explanation
The 8 datasets introduced in Section 3.1 may not always have annotations on all required fields; thus, we modify the first sentence in the Instruction block in Figure 5 with the following variants when appropriate:
  • Only Target is available: ’Provide a brief paragraph to explain step-by-step how the post targets the specified group or entity, and how that leads to the specified Label based on the given Definitions.’
  • Both Target and Rationale are available: ’By elaborating on the provided Annotation, provide a brief paragraph to explain step-by-step how the post targets the specified group or entity, and how that leads to the specified Label based on the given Definitions.’
A.4 Quality Review of Sample Explanations
Elaborating on the criteria outlined in Section 3.1, we review the quality of the generated explanation by GPT-3.5-Turbo on the following items:
  • Grammatically correct
  • Succinct in their justification of the chosen label.
  • Persuasive and logical in their reasoning for the chosen label
We discard explanations that are too verbose, and/or not choosing the label already provided by human annotation (fortunately, this scenario happens rarely, likely due to the presence of existing rationales guiding the extra generated outputs). We also remove explanations that conjure incorrect and/or irrelevant facts to the context to discourage hallucination and encourage a high degree of pertinent to the target post.
Dataset COT-T5-XL Llama 7B Llama 13B
HateCheck 94.5 % 96.5 % 93.5 %
HateXplain 71.0 % 74.0 % 74.0 %
Latent_Hate 50.5 % 51.0 % 44.5 %
Table 5: Percentage out of 200 samples per dataset, where explanations correctly identify at least 1 of the targets listed by human annotators by each model.
A.5 Human Annotation for QA Experiment
A.5.1 QA 1
With approved IRB, we recruit 13 crowdsource workers using the Amazon Mechanical Turk platform to annotate 50 samples per model, for a total of 150 data points for the task described in Section 5.1. The annotators was paid a fair wage at \$15 per hour, and forewarned about the nature of the task. Annotators must be fluent English speakers. We also limit each annotator to no more than 100 posts (60% of the total 150 samples per model) to maintain diversity of opinions. We observe that preference for explanations generated by Pretrained models remains consistent with GPT-4’s.
The demographic breakdown of the 13 annotators are described below:
  • Gender: Female (8), Male (5)
  • Age: 18-29 (2), 30-39 (4), 40-49 (4), 50+ (3)
  • Education: High School (2), 2-year college (5), 4-year college (4), Master’s or Higher (2)
A.5.2 QA 2
For this experiment, each annotation task consists of the explanations generated by 3 models are grouped by the sample. This division resutls in 200 tasks. 6 Amazon Mechanical Turk workers are recruited, with similar qualification criteria as described above.
The demographic breakdown of the 6 annotators are described below:
  • Gender: Female (3), Male (3)
  • Age: 18-29 (1), 30-39 (2), 40-49 (3)
  • Education: 2-year college (2), 4-year college (4)
A.6 QA 3: Target Identification
We investigate the in-domain finetuned models’ capabilities to identify the target of the sentiments expressed by the post. We also randomly select 200 samples from each dataset that have the Target variable annotated by humans, then ask GPT-4 to judge whether the explanations from the models mention at least one of the listed targeted groups using the template in Figure 8. Note that we encourage GPT-4 to consider potential variance of expression and not restrict to exact matches.
Table 5 shows the percentage of accuracy on this task. Similar levels of descending performances are observed in the presented order of test datasets. However, this observation may be an artifact of the differences between the annotations targets among datasets: HateCheck has a limited number of discrete categories while Latent_hate contains multiple combinations of free-text labels. We urge practitioners to consider this factor while curating training data if target identification is desired.
Dataset Definition
Salminen et al.,2018 Neutral : A post that is not offensive to any group of people. Hateful: An offensive post, motivated, in whole or part, by the writer’s bias against an aspect of a group of people.
Qian et al.,2019 Not Hate : Does not contain any attack on people based on protected characteristics. Hate: Contains direct attack on people based on protected characteristics such as race, ethnicity, national origin, religious affiliation, sexual orientation, caste, sex, gender, gender identity, and serious disease or disability.
Sap et al.,2020 Not Offensive : not offensive to anyone. Offensive : denotes the overall rudeness, disrespect, or toxicity of a post. whether a post could be considered offensive to anyone.
Vidgen et al.,2021a Neutral: Content that does not fall into other categories, usually entirely unrelated to abuse, hate, prejudice, or intolerance. Identity Directed Abuse: Content that directs abuse at an identity, which relates to fundamental aspects of individuals’ social position, community and self-representation. An identity includes but is not limited to religion, race, ethnicity, gender, sexuality and sexual preferences, immigration status, nationality, ableness, physical appearance and class. Affiliation Directed Abuse : Content that directs abuse at people who have a voluntary affiliation with a profession, membership, association, ideology, or other well-defined group or collective. Person Directed Abuse : Content that directs abuse at an identifiable person.
Vidgen et al.,2021b Derogation : content which explicitly attacks, demonizes, demeans or insults a group. Animosity : content which expresses abuse against a group in an implicit or subtle manner. Threatening : content which expresses intention to, support for, or encourages inflicting harm on a group, or identified members of the group. Support For Hateful Entities : content which explicitly glorifies, justifies or supports hateful actions, events, organizations, tropes and individuals collectively, entities. Dehumanization : content which perceives or treats people as less than human. Not Hate : content that falls into none of the other categories.
Basile et al.,2019 Hate Speech : language that is used to expresses hatred towards a targeted group or is intended to be derogatory, to humiliate,or to insult the members of the group. may also be language that threatens or incites violence. Offensive Language : may contain offensive terms but targets disadvantaged social groups in a manner that is potentially harmful to them. Neither : language that does not all into either of the other categories.
Hartvigsen et al.,2022 Benign: Text that is not harmful nor offensive to anyone. Toxic: Text that could be seen as harmful to anyone, may contain offensive, rude humor, insults, personal attacks, profanity, aggression, may refer to targeted group with harmful intent that is expressed in stereotypes or lewd manners
Toraman et al.,2022 Hate : target, incite violence against, threaten, or call for physical damage for an individual or a group of people because of some identifying trait or characteristic. Offensive : humiliate, taunt, discriminate, or insult an individual or a group of people in any form, including textual. Normal : does not fall into any of the other categories.
Table 6: Labels and definitions of 8 datasets used to synthesize training corpus. Definitions are lifted directly from their original works, with exceptions of minor adjustment for stylistic consistency.
Dataset Definition
HateCheck Non-hateful: Language that does not contain any abusive sentiment targeted at a protected group. Hateful: Abuse that is targeted at a protected group or at its members being a part of that group. Protected groups include age, disability, gender identity, familial status, pregnancy, race, national or ethnic origins, religion, sex or sexual orientation.
HateXplain Normal : neither hate speech nor offensive. Hate Speech : language which attacks, demeans, offends, threatens, or insults a group based on race, ethnic origin, religion, disability, gender, age, sexual orientation, or other traits. it is not the presence of certain words that makes the text hate speech, rather you should look the context the word is used in the text. Offensive Language : usage of rude, hurtful, derogatory, obscene or insulting language to upset or embarrass people.
Latent_Hate Not Hate : speech or actions that do not involve any form of hatred, prejudice, or discrimination toward individuals or groups based on their characteristics. Explicit Hate: openly expressed, direct forms of hatred and prejudice toward individuals or groups based on their characteristics. Implicit Hate: coded or indirect language that disparages a person or group on the basis of protected characteristics like race, gender, and cultural identity.
Implicit_hate White Grievance : includes frustration over a minority groups perceived privilege and casting majority groups as the real victims of racism. This language is linked to extremist behavior and support for violence. Incitement To Violence : includes flaunting in group unity and power or elevating known hate groups and ideologies. Inferiority Language : implies one group or individual is inferior to another, and it can include dehumanization, denial of a person’s humanity , and toxic language that compares the target with disease, insects, animals . Related to assaults on human dignity, dominance, and declarations of superiority of the in group. Irony : refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. Stereotypes And Misinformation : associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like holocaust denial and other forms of historical negationism. Threatening And Intimidation : conveys a speaker’s commitment to a target’s pain, injury, damage, loss or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation.
Table 7: Labels and definitions for 4 test datasets.
(a) Example from Toraman et al. (2022)
(b) Example from Salminen et al. (2018)
(c) Example from Sap et al. (2020)
Figure 4: Examples drawn from our training corpora showing their native Post, Target and Rationale, along with the corresponding GPT-3.5-Turbo-enhanced explanations. Due to their nature as fragmented annotations, verbatim Rationales are not serviceable explanation, but can serve as guiding signals that leverage GPT’s generative capabilities to construct legible passages with detailed justifications.
Figure 5: Template used to prompt LLM for classification inference.
Figure 6: Template for QA Experiment 1. In this example, EXP_A is preferred.
Figure 7: Template for QA Experiment 2.
Figure 8: Template for QA Experiment 3.